1.     We expect that NA Inc. will pay a dividend of $1.50 per share one year from today, a dividend of $2 per share in years two, and of $3 per share in years three. Assume that the value of the stock at the end of year three to be $18.9. If your required return on NA Inc. stock is 12%, what is the most you would be willing to pay for the stock today if you plan to sell the stock in three years?

 

 


1
Expert's answer
2021-04-15T07:57:50-0400

Answer

"\frac{18.9}{0.12}=157.5"


"\frac{3}{1.5-0.12}=2.17"


"Most willing to pay to the stock=157.5-2.17=155.73"

=$155.73